# A torsion theoretic interpretation for sheaves of modules and Grothendieck topologies on directed categories

### v1

@@ 5.3. Sufficient direction
When ${\mathscr{C}}$ is a noetherian EI category, the following result gives a very transparent description for $J$-torsion $\mathcal{O}$-modules and categories of sheaves of modules. We remark that the topos version (that is, sheaves of sets rather than modules) of the second statement can be deduced from the comparison lemma; see [12, C2.2, Theorem 2.2.3] and [12, C2.2, Example 2.2.4(d)].
@@
(2): Given an $\mathcal{O}$-module $V$, its restriction to ${\mathscr{D}}$ is clearly an $\mathcal{O}_{{\mathscr{D}}}$-module. In this way we obtain an restriction functor $\textsl{res}:\mathcal{O}\operatorname{-Mod}\to\mathcal{O}_{\mathscr{D}}\operatorname{-Mod}$. It has a right adjoint given by the right Kan extension described as follows: given an $\mathcal{O}_{{\mathscr{D}}}$-module $W$ and an object $x\in\operatorname{Ob}({\mathscr{C}})$,
@@ 6. Grothendieck topologies on certain noetherian EI categories
Let $J$ be an exceptional Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ corresponded to a function $\boldsymbol{d}$, and let ${\mathscr{D}}$ be the full subcategory consisting of objects $n$ such that $d_{n}=0$ or $d_{n}=\infty$. Then ${\mathscr{D}}^{{\mathrm{op}}}$ is a $J$-dense subcategory of ${\mathscr{C}}^{{\mathrm{op}}}$ (for a definition of $J$-dense subcategories, see [12, C2.2, Definition 2.2.1]). Therefore, applying the argument in the proof of Theorem 5.16, one has
@@
One can deduce the above results from Lemma 5.10. Indeed, let $J$ be a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Then for each $n\in\mathbb{Z}$, the full subcategory ${\mathscr{C}}_{\geqslant n}$ of objects $m\geqslant n$ is an ideal of ${\mathscr{C}}$ and an EI category of type ${\mathbb{N}}$. These results follows via applying Propositions 6.3, 6.6 and 6.8 to the restriction of $J$ to ${\mathscr{C}}_{\geqslant n}$. When ${\mathscr{C}}$ is the poset $\mathbb{Z}$, they have been proved in [16, Proposition B.31].
7. Applications in group representation theory
@@
In this section we describe some potential applications of our results in group representation theory. A bridge connecting these two areas is the following construction.
@@
Let $G$ be a group, and $\mathcal{H}$ a family of subgroups of $G$. The orbit category ${\mathrm{Orb}}(G,\mathcal{H})$ of $G$ with respect to $\mathcal{H}$ is defined as follows: objects are left cosets $G/H$ with $H\in\mathcal{H}$, and morphisms from $G/H$ to $G/K$ are $G$-equivariant maps. Every such morphism is an epimorphism, and can be represented by an element $g\in G$ such that $gHg^{-1}\subseteq K$. In particular, for every object $G/H$, its endomorphism monoid is a group isomorphic to $N_{G}(H)/H$, so ${\mathrm{Orb}}(G,\mathcal{H})$ is an EI category. For more details, the reader can refer to [17, III.9]. Moreover, it is easy to see that ${\mathrm{Orb}}(G,\mathcal{H})$ is noetherian (resp., artinian) if and only if the poset $\mathcal{H}$ with respect to inclusion is noetherian (resp., artinian).
@@
Let $G$ be a topological group, and let $\mathcal{H}$ be a fundamental system of open subgroups. Impose the atomic Grothendieck topology $J_{a}$ on ${\mathrm{Orb}}(G,\mathcal{H})$, and let $\underline{k}$ be a constant structure sheaf given by a commutative ring $k$. Artin’s theorem [17, III.9 Theorem 1] asserts that
@@
a result described in [24, Subsection 5.3].

### v2

@@ 5.3. Sufficient direction
When ${\mathscr{C}}$ is a noetherian EI category, the following result gives a very transparent description for $J$-torsion $\mathcal{O}$-modules and categories of sheaves of modules.
@@
(2): This is essentially [13, C2.2, Theorem 2.2.3]. For the convenience of the reader, we give a sketched proof. Given an $\mathcal{O}$-module $V$, its restriction to ${\mathscr{D}}$ is clearly an $\mathcal{O}_{{\mathscr{D}}}$-module. In this way we obtain an restriction functor $\textsl{res}:\mathcal{O}\operatorname{-Mod}\to\mathcal{O}_{\mathscr{D}}\operatorname{-Mod}$. It has a right adjoint given by the right Kan extension described as follows: given an $\mathcal{O}_{{\mathscr{D}}}$-module $W$ and an object $x\in\operatorname{Ob}({\mathscr{C}})$,
@@
Remark 5.17.
Note that the above result holds for any rigid Grothendieck topology $J$ on any small category ${\mathscr{C}}^{{\mathrm{op}}}$ by letting ${\mathscr{D}}$ be the full subcategory of ${\mathscr{C}}$ consisting of $J$-irreducible objects; see [13, C2.2, Theorem 2.2.3] and [13, C2.2, Example 2.2.4(d)].
@@ 6. Grothendieck topologies on certain noetherian EI categories
Let $J$ be an exceptional Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ corresponded to a function $\boldsymbol{d}$, and let ${\mathscr{D}}$ be the full subcategory consisting of objects $n$ such that $d_{n}=0$ or $d_{n}=\infty$. Then ${\mathscr{D}}^{{\mathrm{op}}}$ is a $J$-dense subcategory of ${\mathscr{C}}^{{\mathrm{op}}}$ (for a definition of $J$-dense subcategories, see [13, C2.2, Definition 2.2.1]). Therefore, applying the argument in the proof of Theorem 5.16, one has
@@
One can deduce the above results from Lemma 5.10. Indeed, let $J$ be a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Then for each $n\in\mathbb{Z}$, the full subcategory ${\mathscr{C}}_{\geqslant n}$ of objects $m\geqslant n$ is an ideal of ${\mathscr{C}}$ and an EI category of type ${\mathbb{N}}$. These results follows via applying Propositions 6.3, 6.6 and 6.8 to the restriction of $J$ to ${\mathscr{C}}_{\geqslant n}$. When ${\mathscr{C}}$ is the poset $\mathbb{Z}$, they have been proved in [17, Proposition B.31].
7. Applications
@@
In this section we describe some potential applications of our results in representation theory of combinatorial categories and groups.
7.1. Local self-injectivity of infinite full subcategories
Let ${\mathscr{C}}$ be the category of finite sets and injections, or the category of finite dimensional vector spaces over a finite field $\mathbb{F}_{q}$ and $\mathbb{F}_{q}$-linear injections. It has been proved in [11] (see also [22]) that every indecomposable projective ${\mathscr{C}}$-module over a field $k$ of characteristic 0 is also injective. In this subsection we extend this result to their infinite full subcategories.
Recall that an EI category ${\mathscr{C}}$ of type $\mathbb{N}$ is called locally finite if ${\mathscr{C}}(m,n)$ is a finite set for $m,n\in\mathbb{N}$.
Theorem 7.1.
Let ${\mathscr{C}}$ be a locally finite EI category of type $\mathbb{N}$ such that every morphism is monic, ${\mathscr{D}}$ an infinite full subcategory of ${\mathscr{C}}$, and $k$ a field. If every indecomposable projective ${\mathscr{C}}$-module over $k$ is injective, then so is every indecomposable projective ${\mathscr{D}}$-module over $k$.
Proof.
Since ${\mathscr{D}}$ is a full subcategory of ${\mathscr{C}}$, one may impose a rigid Grothendieck topology $J$ on ${\mathscr{C}}^{{\mathrm{op}}}$ such that $J$-irreducible objects in ${\mathscr{C}}$ are precisely objects in ${\mathscr{D}}$. Thus by Remark 5.17, one has
${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J,\,\underline{k})\simeq{\mathscr{D}}\operatorname{-Mod},$
where $\underline{k}$ is the constant structure sheaf on the site $({\mathscr{C}}^{{\mathrm{op}}},J)$, and ${\mathscr{D}}\operatorname{-Mod}$ is the category of ${\mathscr{D}}$-modules over $k$.
Since ${\mathscr{D}}$ is locally finite, to establish the conclusion, it suffices to show that $k{\mathscr{D}}(n,-)$ is injective for each object $n$ in ${\mathscr{D}}$. Note that by [13, C2.2, Theorem 2.2.3], the equivalence from the sheaf category to the module category is given by the restriction functor; and furthermore the restriction of $P(n)=k{\mathscr{C}}(n,-)$ is precisely $k{\mathscr{D}}(n,-)$. Consequently, one only needs to show that $P(n)$ is an injective object in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J,\,\underline{k})$. By Corollary 4.4, this is equivalent to saying that $P(n)$ is a $J$-torsion free injective ${\mathscr{C}}$-module. But we have assumed that $P(n)$ is an injective ${\mathscr{C}}$-module, so it remains to show that $P(n)$ is $J$-torsion free.
Since ${\mathscr{C}}$ is an infinite category, $J$ must be a generic Grothendieck topology by Proposition 6.6. Thus $J(n)\subseteq J_{a}(n)$ for $n\in\mathbb{N}$, where $J_{a}$ is the atomic Grothendieck topology. Accordingly, one has $\mathcal{T}(J)\subseteq\mathcal{T}(J_{a})$ by Lemma 3.6, so every $J_{a}$-torsion free ${\mathscr{C}}$-module over $k$ is also $J$-torsion free. Clearly, $P(n)$ is $J_{a}$-torsion free as every morphism in ${\mathscr{C}}$ is monic, so it is $J$-torsion free as well. This finishes the proof. ∎
7.2. Representations of groups
In this subsection we consider applications of sheaf theory in group representation theory. A bridge connecting these two areas is the following construction.
@@ 7.2. Representations of groups
Let $G$ be a group, and $\mathcal{H}$ a family of subgroups of $G$. The orbit category ${\mathrm{Orb}}(G,\mathcal{H})$ of $G$ with respect to $\mathcal{H}$ is defined as follows: objects are left cosets $G/H$ with $H\in\mathcal{H}$, and morphisms from $G/H$ to $G/K$ are $G$-equivariant maps. Every such morphism is an epimorphism, and can be represented by an element $g\in G$ such that $gHg^{-1}\subseteq K$. In particular, for every object $G/H$, its endomorphism monoid is a group isomorphic to $N_{G}(H)/H$, so ${\mathrm{Orb}}(G,\mathcal{H})$ is an EI category. For more details, the reader can refer to [18, III.9]. Moreover, it is easy to see that ${\mathrm{Orb}}(G,\mathcal{H})$ is noetherian (resp., artinian) if and only if the poset $\mathcal{H}$ with respect to inclusion is noetherian (resp., artinian).
@@
Let $G$ be a topological group, and let $\mathcal{H}$ be a fundamental system of open subgroups. Impose the atomic Grothendieck topology $J_{a}$ on ${\mathrm{Orb}}(G,\mathcal{H})$, and let $\underline{k}$ be a constant structure sheaf given by a commutative ring $k$. Artin’s theorem [18, III.9 Theorem 1] asserts that
@@
a result described in [26, Subsection 5.3].